Beccaaaa:

Help with math please :> Parallel to y=-1/5x-7 through (2,3)

DeadOfWinter:

@astrid1

DeadOfWinter:

@extrinix

SatanBunneh:

@stopthataxel

SatanBunneh:

Answer.

Swirxy:

uhhhhhhhhhhhhh

Swirxy:

idek that

Beccaaaa:

It's all good thanks tho.

Swirxy:

k

Swirxy:

yw

SatanicDevil:

@joe348

TooCoolNot:

Idk what this means but he Pmed me dis: StopThatAxel a minute ago All lines that are parallel to y=−15x+7 y = - 1 5 x + 7 have the same slope of −15 - 1 5 . mparallel=−15 m parallel = - 1 5. Step 4. Enter a problem.

Beccaaaa:

@toocoolnot wrote:
Idk what this means but he Pmed me dis: StopThatAxel a minute ago All lines that are parallel to y=−15x+7 y = - 1 5 x + 7 have the same slope of −15 - 1 5 . mparallel=−15 m parallel = - 1 5. Step 4. Enter a problem.
Ngl, that's confusing..

TooCoolNot:

@beccaaaa wrote:
@toocoolnot wrote:
Idk what this means but he Pmed me dis: StopThatAxel a minute ago All lines that are parallel to y=−15x+7 y = - 1 5 x + 7 have the same slope of −15 - 1 5 . mparallel=−15 m parallel = - 1 5. Step 4. Enter a problem.
Ngl, that's confusing..
Well.. that's the only way to put it cause we can't give the actual answer ;-;

Beccaaaa:

True, i'll figure something out. Tyy

xhikari:

so the line we're making must pass through (2,3) .. if we graph that we can see that the y-intersect is around 4.5 so then your equation for the parallel line would be y= -1/5 + 4.5 because the slopes are the same since the lines are parallel, does this help?

Beccaaaa:

@xhikari wrote:
so the line we're making must pass through (2,3) .. if we graph that we can see that the y-intersect is around 4.5 so then your equation for the parallel line would be y= -1/5 + 4.5 because the slopes are the same since the lines are parallel, does this help?
A little. Thank you so much :>

xhikari:

@beccaaaa wrote:
@xhikari wrote:
so the line we're making must pass through (2,3) .. if we graph that we can see that the y-intersect is around 4.5 so then your equation for the parallel line would be y= -1/5 + 4.5 because the slopes are the same since the lines are parallel, does this help?
A little. Thank you so much :>
no problem :3

fro28rub:

uhhh.... u must be in highschool, im in middle school 7th grade lol idk

jaquan456:

Ok Beccaaaa I don't have a clue I learned this but I know a app That's totally free that gives you the answer right away....MathWay -JaQuan

Beccaaaa:

Thank you

jaquan456:

Your welcome lemme know if it didn't work ill do it for you!

Forku:

ik this was but ima answer anyway 1/5= 5 then 5 X -7= -35

Swirxy:

@forku wrote:
ik this was but ima answer anyway 1/5= 5 then 5 X -7= -35
this was a month ago

Swirxy:

@swirxy wrote:
@forku wrote:
ik this was but ima answer anyway 1/5= 5 then 5 X -7= -35
this was a month ago
there is no need to reply

Swirxy:

All lines that are parallel to y=−15x+7 y = - 1 5 x + 7 have the same slope of −15 - 1 5 . m parallel=−15 m parallel = - 1 5. Step 4. image of graph.

DeadOfWinter:

@swirxy wrote:
All lines that are parallel to y=−15x+7 y = - 1 5 x + 7 have the same slope of −15 - 1 5 . m parallel=−15 m parallel = - 1 5. Step 4. image of graph.
Bruh

Swirxy:

@deadofwinter wrote:
@swirxy wrote:
All lines that are parallel to y=−15x+7 y = - 1 5 x + 7 have the same slope of −15 - 1 5 . m parallel=−15 m parallel = - 1 5. Step 4. image of graph.
Bruh
wut.

DeadOfWinter:

Nun